OpenStudy (anonymous):

How many committees of six can be formed out of a group of twelve Republicans and ten Democrats if no more than four are Democrats?

OpenStudy (valpey):

First count the possible D/R splits: 4D, 2R 3D, 3R 2D, 4R 1D, 5R 0D, 6R Then for each of these x D, y R, count how many ways you can pick x out of 10 times the number of ways you can pick y out of 12.

OpenStudy (kropot72):

The number of committees is given by: \[12C6+(10\times 12C5)+(10C2\times 12C4)+(10C3\times 12C3)+(10C4\times 12C2)\]

OpenStudy (valpey):

kropot72 is correct. Do you understand combinatorial notation?